OpenStudy (anonymous):

Write an equation of the line that contains the median of triangle RST from S to line RT. S(-4,7), R(0,1), T(4,7).....HELP!!!!!!!

OpenStudy (anonymous):

The median goes from one angle to the midpoint of the opposite side. The midpoint of RT is (2, 4). So the median goes through (2, 4) and (-4, 7). The slope of the line is (7 - 4)/(-4 -2) or -1/2. We can use the point/slope form: y-4 = -1/2(x-2). In slope/intercept form, this is y = -1/2x + 5.